OpenStudy (anonymous):

(x + 3)^2/3 + 2(x + 3)^1/3 - 15 = 0

OpenStudy (anonymous):

try solving \[z^2+2z-15=0\] and then replace z by \[\sqrt[3]{x+3}\]

OpenStudy (phi):

rename (x+3)^(1/3) as y so you have y^2 + 2y -15=0 solve for y x= y^(-1/3)-3